OpenStudy (anonymous):

you want to buy a new bike . you have $50 in your saving account. you earn $12 for every lawn that you mow . use the equation B=50+12, where b is the total cost of the bike and m is the number of the lawn that you mow . if the total cost of the bike is $165 how many lawns will you mow if you want to purchase the bike ( show work) :)

OpenStudy (anonymous):

It should be B=50+12m, right? 165=50+12m 165-50=12m 115=12m m=115/12 m=9.5833333333 so, 10 lawns.